Suppose the risk-free rate is 2.03% and an analyst assumes a market risk premium of 5.18%. Firm A just paid a dividend of $1.2 decimal places
Answer #1

Value of Firm B=Number of shares*D0*(1+g)/(rf+b*MRP-g)=198*1.72*1.022/(2.03%+0.71*5.18%-2.20%)=9922.23958 million